OpenStudy (anonymous):

write the line that passes through (-1,5) with aslope of 2 in standard form

OpenStudy (anonymous):

slope = \[\frac{ y -y _{1}}{ x-x _{1} }\] where slope = 2, \[y _{1} = 5\] and so\[y-5 = 2(x+1)\] and so y = 2x +7

sam (.sam.):

Use the point-slope form \(y-y_1=m(x-x_1)\) Where \((x_1,y_1)\) and slope = m

OpenStudy (anonymous):

oh okay